# Data Distribution Bottlenecks in Grounding Language Models to Knowledge Bases

### 30 Jan 2024 v2

@@ Abstract
Language models (LMs) have already demonstrated remarkable abilities in understanding and generating both natural language and formal language. Despite these advances, their integration with real-world environments such as large-scale knowledge bases (KBs) remains an underdeveloped area, affecting applications such as semantic parsing and indulging in "hallucinated" information. This paper is an experimental investigation aimed at uncovering the robustness challenges that LMs encounter when tasked with knowledge base question-answering (KBQA). The investigation covers scenarios with inconsistent data distribution between training and inference, such as generalization to unseen domains, adaptation to various language variations, and transferability across different datasets. Our comprehensive experiments reveal that even when employed with our proposed data augmentation techniques, advanced small and large language models exhibit poor performance in various dimensions. While the LM is a promising technology, the robustness of the current form in dealing with complex environments is fragile and of limited practicality because of the data distribution issue. This calls for future research on data collection and LM learning paradigms [^1] Code and data will be public upon acceptance..
Data Distribution Bottlenecks in Grounding Language Models to Knowledge Bases
Yiheng Shuthanks: * Work performed during internship at Microsoft Research Asia. The Ohio State University shu.251@osu.edu                        Zhiwei Yu Microsoft zhiwyu@microsoft.com
@@ 1 Introduction
Language models (LMs), such as BERT Devlin et al. (2019), T5 Raffel et al. (2020), and the GPT series Ouyang et al. (2022); OpenAI (2023), have demonstrated impressive capabilities in understanding and generating natural or formal language, highlighting the potential for artificial general intelligence (AGI). However, several obstacles must be overcome to achieve this goal. For example, to mitigate the issue of "hallucinated" information and expand the range of LM applications, researchers have focused on grounding LMs to real-world environments (e.g., Web, database, knowledge base) Nakano et al. (2021); Menick et al. (2022), which enables real-time fact-checking, data validation and thereby improve the reliability of model responses. Knowledge base (KB) is one underdeveloped environment among them. The task of Knowledge Base Question Answering (KBQA) aims to parse natural language queries using KBs, such as Freebase Bollacker et al. (2008), and Wikidata Vrandecic and Krötzsch (2014). Now, numerous LM-driven models Das et al. (2021); Hu et al. (2022) continue to achieve higher scores on KBQA benchmarks, but grounding LMs to KBs has not been fully examined for robustness and reliability. A few critical gaps in existing works prompt the necessity for this work. First, while the evaluation of LMs typically occurs in natural language tasks, the challenge escalates when grounding these models to real-world environments like KBs Liu et al. (2023), where the data contains structured data instead of all unstructured natural language. Second, the metrics of the KBQA benchmark are often shallow and the robustness of the model is not adequately evaluated. Finally, recent surveys on KBQA Lan et al. (2022); Gu et al. (2022b) have largely omitted the strides made in the development and application of LMs, especially large language models (LLMs). As a result, there remains a clear gap in understanding the robustness challenges and solutions specific to grounding LMs to KBs.
Figure 1: Inconsistent KBQA data distribution between training and inference.
@@
For deep learning models, robustness is closely related to data distribution Hendrycks et al. (2020). The efficacy of language models (LMs) is not just a product of their architecture, but also a reflection of the data on which they are trained. In simpler tasks and well-defined domains, large-scale corpora have been collected and used for effective training Touvron et al. (2023). However, real-world environments are rarely so accommodating, e.g., large KBs contain complex structures and schema items, and building a representative corpus is more challenging. Inconsistency in data distribution during training and inference, as shown in Figure 1, may negatively impact the performance and robustness of LMs. Given this backdrop, it becomes essential to approach the task of grounding LMs from a multifaceted standpoint. Specifically, we believe that a detailed understanding of this grounding task requires exploring environmental, linguistic, and model learning aspects.
Through these explorations, this paper aims to provide a more holistic understanding of the challenges and opportunities that arise in the LM grounding and KBQA benchmarking processes. We review existing works and identify several challenges, including: 1) generalization to unseen domains at the schema level Gu et al. (2021), 2) adaptation to paraphrases featuring diverse language variations Su et al. (2016), 3) transferability across datasets with novel schema items, query patterns, and linguistic styles Cao et al. (2022b), and 4) few-shot in-context learning capabilities of grounding LLMs Li et al. (2023). These aspects allow us to dissect and critique the robustness and applicability of LMs when operating in complex real-world scenarios, thereby fulfilling our objective of effective interaction between LMs and environments.
@@
To substantiate the presence of these challenges and obtain additional empirical insights, we undertake a comprehensive experimental investigation for the above aspects. To mitigate the negative effects of inconsistent data distributions in our experiments, we propose a data augmentation method for any LM and a retrieval augmentation method for any LLM. Our findings reveal that even when employed with such techniques (the highest EM score is achieved on the GrailQA benchmark), advanced small and large LMs still fall short of effectively tackling the majority of these challenges. A striking example is the large difference between the best practice without WebQSP Yih et al. (2016) fine-tuning (F1 43.0%) compared to the fine-tuned state-of-the-art (F1 79.6%) (§6), suggesting the weak robustness of LM-driven KBQA models on an unseen dataset. Such observations highlight an urgent need for future research in data collection methodologies and LM learning paradigms. Meanwhile, we expect the evaluation approaches in this paper to provide a reference for future benchmark construction, developing metrics that take robustness into account (§3).
2 Preliminary
In this paper, the knowledge base (KB) refers to an RDF [^2] Resource Description Framework by W3C standard. graph, consisting of triples $(s,r,o)$, where $s$ is a subject, $r$ is a relation and $o$ is an object. Logical form refers to a formal query that can be executed on a KB like SPARQL and S-expression Gu et al. (2021). Schema refers to the rdfs:Class (class) and rdf:Property (relation) in the RDF Schema [^3] https://www.w3.org/TR/rdf12-schema/.
Benchmark KB Size LF Generalization Para.
WebQuestions Berant et al. (2013) Freebase 5,810 N/A i.i.d.
SimpleQuestions Bordes et al. (2015) Freebase 108,442 N/A i.i.d.
WebQuestionsSP Yih et al. (2016) Freebase 4,737 SPARQL i.i.d.
GraphQuestions Su et al. (2016) Freebase 5,166 Graph query comp.+zero
LC-QuAD Trivedi et al. (2017) DBpedia 5,000 SPARQL i.i.d.
CWQ Talmor and Berant (2018) Freebase 34,689 SPARQL i.i.d.
LC-QuAD 2.0 Dubey et al. (2019) Wikidata 30,000 SPARQL i.i.d.
SQB Wu et al. (2019) Freebase 108,443 N/A i.i.d.+zero
CFQ Keysers et al. (2020) Freebase 239,357 SPARQL comp
GrailQA Gu et al. (2021) Freebase 64,331 S-expression i.i.d.+comp.+zero
KQA Pro Cao et al. (2022a) Wikidata 117,970 KoPL i.i.d.
QALD series Perevalov et al. (2022) DBpedia 558 SPARQL comp.
@@ 2 Preliminary
Table 1: Selected KBQA benchmarks. LF: logical forms. Generalization settings follow Gu et al. (2021). i.i.d. denotes that the schema distribution in the test set is the same as the training set. comp. and zero denote compositional and zero-shot generalization, respectively. Para. denotes paraphrases, i.e., questions containing the same semantics (machine-generated paraphrases are not included).
3 Robustness Challenge
@@
In this paper, the robustness of a KBQA model refers to its ability to adapt to various natural language inputs and maintain consistent performance when data distribution shifts. Due to the data discrepancy between the training corpus of LMs and KB environments, LMs face challenges from environmental, linguistic, and model learning aspects.
@@ 3.1 Environmental Aspect
One of the principal challenges from an environmental perspective is schema-level generalization. The RDF Schema offers a data-modeling vocabulary, which is essential for querying a KB. As shown in Table 1, most KBQA benchmarks operate under the assumption that the schema distribution is identically and independently distributed (i.i.d.) between training and testing scenarios. This means that the schema items encountered during testing are already familiar to the model from the training data. However, this assumption often falls short in large-scale KBs, which contains thousands of schema items. Currently, only a limited number of benchmarks aim to evaluate the ability to generalize at the schema level, especially when it comes to handling non-i.i.d. schema items. The SimpleQuestions-Balance dataset Wu et al. (2019) reconfigures the i.i.d. SimpleQuestions dataset Petrochuk and Zettlemoyer (2018) to ensure that half of the questions in the testing (or validation) set are mapped to KB relations not present in the training data. GrailQA Gu et al. (2021) is structured to evaluate three levels of schema-level generalization: i.i.d. (25%), compositional generalization (25%), and zero-shot generalization (50%). GraphQuestions Su et al. (2016) presents an even more rigorous test as its testing set features schema items seldom encountered in the training data. Given the impracticality of maintaining the i.i.d. assumption in real-world applications, these datasets offer a more realistic portrayal of the generalization challenges that QA systems face in practical settings. Despite recent advances Shu et al. (2022); Gu et al. (2022a), challenges in compositional and zero-shot generalization are far from solved.
@@ 3.2 Linguistic Aspect
Natural language is variable in form, making question understanding challenging for KBQA models. One common way this variety shows up is through paraphrasing. In this paper, a paraphrase set denotes different ways to express the same logical form, as illustrated in Table 14. To gauge how well a KBQA model can adapt to these different forms of natural language, a straightforward approach is to test if the model can accurately answer paraphrased questions that it has already answered correctly before. Unfortunately, as shown in Table 1, many KBQA benchmarks do not account for paraphrasing with only one utterance for each logical form. Exceptionally, some datasets Su et al. (2016); Dubey et al. (2019); Gu et al. (2021) are based on automatically generated logical forms and include multiple natural language expressions for the same logical form (template). These data characteristics highlight the difficulties in adapting to paraphrased questions.
@@ 3.3 Integrated Aspect
Evaluating KBQA benchmarks often hinges on a single dataset, thereby complicating the task of ascertaining the model performance consistency across novel scenarios. This form of robustness, termed as cross-dataset transfer in this paper, combines both the environmental and linguistic aspects discussed earlier and is more difficult to achieve. This is because construction methods vary across datasets, as do schema distributions and natural language expressions. Specifically, KBQA dataset construction generally falls into two distinct categories: 1) Graph Search and Crowdsourcing: in this approach, logical forms or triples are initially extracted from a KB, where structures or operators of logical form are usually finite. Subsequently, these are converted into natural language utterances through crowdsourcing techniques Bordes et al. (2015); Trivedi et al. (2017). 2) Human Curation and Parsing: logical forms are labeled directly from human-provided utterances Berant et al. (2013); Perevalov et al. (2022). Existing works Gu et al. (2021); Cao et al. (2022b) suggest that models pre-trained on large-scale datasets can adapt reasonably well to other target datasets, such as WebQuestionsSP Yih et al. (2016). However, the necessity for fine-tuning these pre-trained models on the intended target dataset remains imperative for achieving optimal performance. That is, despite the advantages offered by pre-training on expansive KBQA datasets, models still encounter challenges in transferring directly to previously unseen target datasets while sustaining high performance.
3.4 Learning Aspect
@@
Aside from considering environmental and linguistic factors, it is crucial to focus on the model learning process. Recently, LLMs like GPT series OpenAI (2023) have demonstrated exceptional capabilities across a variety of tasks, outperforming smaller yet potent LMs such as BERT Devlin et al. (2019) and T5 Raffel et al. (2020). Despite these advancements, these LLMs face substantial challenges when interacting with environments. One notable issue is their predominant reliance on an in-context learning paradigm as opposed to fine-tuning, as a trade-off between computational cost and model efficiency. In comparison to fine-tuning, in-context learning offers the advantage of reduced training costs but at the expense of the lacking perception of the environment. Inconsistent data distribution between natural language pre-training and reasoning over structured knowledge contexts leads to poor performance. For instance, a discernible performance gap exists between KBQA models that employ in-context learning with Codex Chen et al. (2021a) and those built on fine-tuned LMs Gu et al. (2022a); Li et al. (2023). Therefore, it is crucial to consider the limitations of the commonly used in-context learning paradigm and improve the grounding approaches.
4 Approach
The goal of our proposed approaches is to reduce the negative effects of inconsistent data distributions and to maximize the potential of LMs in our experimental investigations. We introduce both data augmentation and retrieval augmentation techniques.
4.1 Data Augmentation for LMs
Off-the-shelf datasets of limited size may make LM easily overfitted and not adaptable to large KBs. To address the problem that many domains in the KB are often not collected as training data, we propose a data augmentation method named Graph seArch and questIon generatioN (GAIN). GAIN applies to KBQA corresponding to logical forms or triples, and scales data volume and distribution through four steps: 1) Graph search: Sampling logical forms or triples from arbitrary domains in the KB, without being restricted to any particular KBQA dataset. 2) Training question generator on existing KBQA datasets, i.e., learning to convert logical forms or triples into natural language questions. 3) Verbalization: Using the question generator from step 2 to verbalize sampled logical forms or triples from step 1, thus creating synthetic questions. 4) Training data expansion: Before fine-tuning any neural models on KBQA datasets, GAIN-synthetic data can be used to train these models or to expand the corpus of in-context samples for LLMs. That is, as a data augmentation method, GAIN is not a KBQA model, but it is used to augment a base KBQA model.
4.2 Retrieval Augmentation for LLMs
As the trade-off between cost and effectiveness, we experiment with the prevalent in-context learning paradigm but attempt to improve the quality of in-context samples. We use advanced retrieval methods based on smaller LMs as plug-ins to augment the LLM, similar to the SuperICL approach Xu et al. (2023). Specifically, our steps to generate an LLM prompt for each question include the following. 1) Given an input question, we retrieve $k$ questions ($k$-shot) with BM25 Robertson et al. (2009) from the corpus (the combination of KBQA training set and the GAIN-synthetic dataset). 2) The role of retrieval augmentation for KB environments has been shown by fine-tuned LMs Shu et al. (2022). To assist with grounding LLM, we retrieve KB contexts with off-the-shelf retrievers for $k$ samples and the input question [^4] The prompt example is demonstrated in Appendix A..

### 09 Feb 2024 v3

# Distribution Shifts Are Bottlenecks: Extensive Evaluation for Grounding Language Models to Knowledge Bases

@@ Abstract
Grounding language models (LMs) to knowledge bases (KBs) helps to obtain rich and accurate facts. However, it remains challenging because of the enormous size, complex structure, and partial observability of KBs. One reason is that current benchmarks fail to reflect robustness challenges and fairly evaluate models. This paper analyzes whether these robustness challenges arise from distribution shifts, including environmental, linguistic, and modal aspects. This affects the ability of LMs to cope with unseen schema, adapt to language variations, and perform few-shot learning. Thus, the paper proposes extensive evaluation protocols and conducts experiments to demonstrate that, despite utilizing our proposed data augmentation method, both advanced small and large language models exhibit poor robustness in these aspects. We conclude that current LMs are too fragile to navigate in complex environments due to distribution shifts. This underscores the need for future research focusing on data collection, evaluation protocols, and learning paradigms. [^1] Code and data are available at https://github.com/yhshu/Distribution-Shifts-for-KBQA.
Distribution Shifts Are Bottlenecks: Extensive Evaluation for
Grounding Language Models to Knowledge Bases
Yiheng Shuthanks: * Work performed when the author was graduate student at Nanjing University. The Ohio State University shu.251@osu.edu                        Zhiwei Yu Microsoft zhiwyu@microsoft.com
@@ 1 Introduction
Language models (LMs), such as BERT Devlin et al. (2019), T5 Raffel et al. (2020), and the GPT series Ouyang et al. (2022); OpenAI (2023), have demonstrated impressive capabilities in understanding and generating languages, highlighting the potential for artificial general intelligence (AGI). However, a major obstacle to achieving this goal is that LMs mainly built on natural languages are not yet well-grounded to real-world environments, such as knowledge base (KB), an environment of enormous size, complex structure, and only partially observable to LM.
Figure 1: Distribution shifts during grounding in the case of KBQA. Training and inference using LM face completely different data distributions. We call for these shifts to be reflected in the evaluation protocols.
Though LMs are highly skilled at natural language question answering (QA) today, the task of Knowledge Base Question Answering (KBQA) aims to parse natural language queries into formal queries on KBs, such as Freebase Bollacker et al. (2008) and Wikidata Vrandecic and Krötzsch (2014). The significance of this task lies in building language agents on complex environments Su (2023), rather than merely recalling answers from the LM’s stored knowledge.
Now, numerous LM-driven models Das et al. (2021); Hu et al. (2022) continue to achieve higher F1/Hits@1 scores on KBQA benchmarks. However, achieving higher scores does not necessarily guarantee the development of robust and dependable models. We still need to ask whether such improvements apply to extensive scenarios, as benchmarks almost always create questions via crowdsourcing and evaluate with simplistic metrics (Table 1). These benchmarks may not fully represent the diverse scenarios encountered in real-world applications, which raises concerns about the robustness of LM-driven models. Thus, our research aims to bridge this gap by exploring the limitations of current KBQA benchmarks and proposing more comprehensive evaluation protocols.
@@
To achieve this goal, we need to grasp the key factors in robustness. For modern deep learning systems, the amount of training data could be extremely rich, but robustness is closely related to data distribution Hendrycks et al. (2020). In the general area of natural language processing (NLP), large-scale corpora have been collected and used for effective training Touvron et al. (2023). However, real-world environments are rarely so accommodating, e.g., large KBs contain complex structures and schema items, and building a large-scale and representative corpus is quite challenging. The problem of inconsistent data distribution during training and inference, i.e., distribution shifts as shown in Figure 1, may negatively impact the performance and robustness of grounded LMs.
This paper analyzes the distribution shifts from multiple aspects to understand both KBQA benchmarking and LM grounding better. We review existing works and identify several challenges. 1) Environmental aspect: generalization to unseen domains at the schema [^2] Schema denotes rdfs:Class (class) and rdf:Property (relation) here. level Gu et al. (2021); 2) Linguistic aspect: adaptation to paraphrases featuring diverse language variations Su et al. (2016); 3) Integrated aspect: transferability across datasets with both novel schema items and linguistic styles Cao et al. (2022b); 4) Modal aspect: few-shot learning ability of large language models (LLMs) for the unfamiliar KB modality Li et al. (2023). These challenges guide us to build more rigorous evaluation protocols for multiple aspects (§4).
@@
To measure the impact of distribution shifts, we conduct extensive experiments under the proposed evaluation protocols. To present a fair evaluation under such rigorous protocols, we implement a data augmentation method for evaluated LMs and a retrieval augmentation method for evaluated LLMs (§5.1). Our findings reveal that even when employed with such methods and the highest EM scores are achieved on the GrailQA benchmark Gu et al. (2021), advanced small and large LMs still fall short of effectively tackling the majority of these challenges. A striking example is the large difference between the best practice without WebQSP Yih et al. (2016) fine-tuning (F1 43.0%) compared to the fine-tuned state-of-the-art (F1 79.6%) (§6), suggesting the weak robustness of LM-driven KBQA models on an unseen dataset. Such negative observations highlight an urgent need for future research in data collection methodologies and LM learning paradigms. Meanwhile, we expect our evaluation protocols to provide a reference for future benchmark construction, developing metrics that consider robustness.
Our contributions include 1) A systematic analysis of the robustness challenges of grounded LMs and a strong advocacy of a multi-aspect evaluation protocol for KBQA. 2) Extensive experiments quantitatively unveil the existence of these challenges and the vulnerability of both small and large LMs. 3) Insights into improving the robustness of grounded LMs, including data collection and learning paradigms.
2 Related Work
Compared to existing research, the necessity of this study is threefold. First, while LMs are commonly evaluated on natural language tasks Hupkes et al. (2022); Patel et al. (2022), the complexity increases significantly when these models are applied to environments like KBs, where data is structured rather than purely unstructured natural language Liu et al. (2023). Second, the evaluation protocols used in KBQA benchmarks tend to be uniform, leading to an insufficient evaluation of model robustness Gu et al. (2021). Finally, recent reviews in this field Lan et al. (2022); Gu et al. (2022b) have largely overlooked the advancements in the development and application of LMs, particularly LLMs.
Benchmark KB Size LF Generalization Paraphrases Metrics
WebQuestions Berant et al. (2013) Freebase 5,810 N/A i.i.d. F1
SimpleQuestions Bordes et al. (2015) Freebase 108,442 N/A i.i.d. Acc
WebQuestionsSP Yih et al. (2016) Freebase 4,737 SPARQL i.i.d. F1
GraphQuestions Su et al. (2016) Freebase 5,166 Graph query comp.+zero F1
LC-QuAD Trivedi et al. (2017) DBpedia 5,000 SPARQL i.i.d. F1
CWQ Talmor and Berant (2018) Freebase 34,689 SPARQL i.i.d. Hits
LC-QuAD 2.0 Dubey et al. (2019) Wikidata 30,000 SPARQL i.i.d. F1
SQB Wu et al. (2019) Freebase 108,443 N/A i.i.d.+zero Acc
CFQ Keysers et al. (2020) Freebase 239,357 SPARQL comp Acc
GrailQA Gu et al. (2021) Freebase 64,331 S-expression i.i.d.+comp.+zero EM, F1
KQA Pro Cao et al. (2022a) Wikidata 117,970 KoPL i.i.d. Acc
QALD series Perevalov et al. (2022) DBpedia 558 SPARQL comp. F1
@@ 2 Related Work
Table 1: Selected KBQA benchmarks. LF denotes logical forms. Generalization settings follow Gu et al. (2021). i.i.d. denotes that the schema distribution in the test set is the same as the training set. comp. and zero denote compositional and zero-shot generalization, respectively. Paraphrases are questions containing the same semantics (machine-generated paraphrases are not included). Acc denotes accuracy. EM denotes exact match.
3 Challenges from Distribution Shifts
@@
In this paper, the robustness of a KBQA model refers to its ability to adapt to various natural language inputs and maintain consistent performance when data distribution shifts. Due to the distribution shifts between the training corpus of LMs and KB environments, grounding LMs face robustness challenges from environmental, linguistic, and modal aspects.
@@ 3.1 Environmental Aspect
A primary environmental challenge is schema-level generalization. The RDF Schema provides a data-modeling vocabulary crucial for querying a KB. Table 1 indicates most KBQA benchmarks assume a consistent schema distribution between training and testing. However, this often does not hold in large KBs with numerous schema items. Few benchmarks, like the reconfigured SimpleQuestions-Balance dataset Wu et al. (2019) and GrailQA Gu et al. (2021), address non-i.i.d. schema items and varying levels of schema-level generalization. GraphQuestions Su et al. (2016) provides a stringent test with seldom-seen schema items in training. Given real-world non-i.i.d. complexities, these datasets better represent the practical generalization challenges. Yet, despite progress Shu et al. (2022); Gu et al. (2022a), compositional and zero-shot generalization are far from solved. We stress that this challenge applies to the overall KBQA semantic parsing process, especially to important subtasks such as relation linking.
@@ 3.2 Linguistic Aspect
Various natural language expressions make question understanding challenging for KBQA models. One common way this variety shows up is through paraphrasing. In this paper, a paraphrase set denotes different ways to express the same logical form, as illustrated in Table 14. Paraphrase adaptation is an intuitive form of the ability to comprehend variable language expressions. It could be measured by whether a model is able to accurately answer paraphrased questions that this model has already answered correctly before. Unfortunately, as shown in Table 1, many KBQA benchmarks do not account for paraphrasing with only one utterance for each logical form. Exceptionally, some datasets Su et al. (2016); Dubey et al. (2019); Gu et al. (2021) are based on automatically generated logical forms and include multiple natural language expressions for the same logical form (template). These data characteristics highlight the difficulties in adapting to paraphrased questions.
@@ 3.3 Integrated Aspect
Evaluating KBQA benchmarks often hinges on a single dataset, thereby complicating the task of ascertaining the model performance consistency across novel scenarios. This form of robustness, termed as cross-dataset transfer in this paper, combines both the environmental and linguistic aspects discussed earlier and is more difficult to achieve. This is because construction methods vary across datasets, as do schema distributions and natural language expressions. Specifically, KBQA dataset construction generally falls into two distinct categories: 1) Graph Search and Crowdsourcing: in this approach, logical forms or triples are initially extracted from a KB, where structures or operators of logical form are usually finite. Subsequently, they are converted into natural language utterances through crowdsourcing methods Bordes et al. (2015); Trivedi et al. (2017). 2) Human Curation and Parsing: logical forms are labeled directly from human-provided utterances Berant et al. (2013); Perevalov et al. (2022). Existing works Gu et al. (2021); Cao et al. (2022b) suggest that models pre-trained on large-scale datasets can adapt reasonably well to other target datasets, such as WebQSP Yih et al. (2016). However, the necessity for fine-tuning these pre-trained models on the intended target dataset remains imperative for achieving optimal performance. Despite the advantages offered by pre-training on expansive KBQA datasets, models still encounter challenges in transferring directly to previously unseen target datasets while sustaining high performance.
3.4 Modal Aspect
@@
Aside from considering environmental and linguistic factors, focusing on the modal aspect is also crucial. Recently, LLMs like GPT series OpenAI (2023) have demonstrated exceptional capabilities across a variety of tasks, outperforming smaller yet potent LMs such as BERT Devlin et al. (2019) and T5 Raffel et al. (2020). Despite these advancements, these LLMs face substantial challenges when interacting with environments. One notable issue is they predominantly rely on an in-context learning paradigm as opposed to fine-tuning, as a trade-off between computational cost and model efficiency. In comparison to fine-tuning, in-context learning offers the advantage of reduced training costs but at the expense of being forced to reason over the unfamiliar modality. Distribution shifts between natural language pre-training and reasoning over structured knowledge contexts could lead to poor performance. For instance, a discernible performance gap exists between KBQA models that employ in-context learning with Codex Chen et al. (2021a) and those built on fine-tuned LMs Gu et al. (2022a); Li et al. (2023). However, the empirical specifics of this difference are not yet clear, leaving us with an inadequate understanding of the limitations of in-context learning and ways to improve grounding with LLMs.
4 Evaluation Protocols
Regarding these challenges, we introduce extensive protocols for evaluating LMs in several aspects overlooked by current benchmarks.
4.1 Evaluating Environmental Aspect
To set the environmental schema-level generalization scenario, we use GrailQA Gu et al. (2021) and GraphQuestions Su et al. (2016) datasets. GrailQA contains three generalization levels: i.i.d. (25%), compositional (25%), and zero-shot (50%). GraphQuestions has no seen relations in the test set. We also use SimpleQuestions-Balance (SQB) Wu et al. (2019) for the relation linking task (an important KBQA sub-task), where 50% of the samples contain unseen relations.
4.2 Evaluating Linguistic Aspect
To set a paraphrase adaptation scenario, we use GrailQA Gu et al. (2021) and GraphQuestions Su et al. (2016) datasets. To evaluate adaptability to paraphrases (§3.2), we propose a new metric, the standard deviation (std) of EM/F1 for questions of each logical form template. As shown in Equation 1, suppose there are $n$ sets of paraphrases in the dataset, each set of paraphrases corresponds to a logical form template with $m$ natural language expressions, and the F1 score obtained by the KBQA model on the $j$-th question of the $i$-th set of paraphrases is $F1_{i,j}$. The metric $Std_{F1}$ first calculates the standard deviation of the F1 scores obtained by the model on the $m$ questions for each set of paraphrases and then calculates the average of the $n$ standard deviations. This metric is used to measure the robustness of the model to different representations of the same semantics, i.e., whether it can cope with diverse natural language expressions. A lower standard deviation indicates that the model is more adaptive to different expressions. $Std_{EM}$ is calculated in the same way.
4.3 Evaluating Integrated Aspect
To emulate a real-world scenario with unknown schema and linguistic distribution for the integrated aspect, we evaluate the performance of pre-trained models on the unseen human-curated WebQSP Yih et al. (2016) dataset, where the questions are derived from search logs and are more realistic. This is a scenario where the distribution changes significantly, as most benchmarks create questions by sampling logical forms and annotating natural language questions via crowdsourcing Lan et al. (2022), where questions are confined to sampled logical forms.
4.4 Evaluating Modal Aspect
To test the capability of the in-context learning paradigm to inference from the KB modality rather than from pure texts, we retrieve structured KB contexts as prompt to evaluate the LLM without particular fine-tuning on KB (§5.1.2). KBs are structured and expansive, but in this case, LLM can only encode a portion of linearized KB contexts.